The perimeter of a quarter circle is 7.14 kilometers. What is the quarter circle's radius?


1
Expert's answer
2022-06-12T17:13:13-0400

"Perimeter ~of ~a~ circle= 2\pi r"


"Perimeter ~of ~a~ quarter ~circle= \cfrac{2\pi r}{4}= \cfrac {πr}{2}"



"Perimeter~of~a~quarter~circle= 7.14"


"\cfrac {πr}{2}= 7.14"



"\u03c0r = 7.14*2 = 14.28"


"r= \cfrac {14.28}{π}= \cfrac {14.28}{3.14}= 4.548"



Answer: 4,548 km